# Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization

### v2

@@ Abstract
Realistic user simulation is crucial for training and evaluating task-oriented dialogue (TOD) systems, yet creating simulators that accurately replicate human behavior remains challenging. A key property of effective simulators is their ability to expose failure modes of the systems they evaluate. We present an adversarial training framework that iteratively improves user simulator realism through a competitive dynamic between a generator (user simulator) and a discriminator. Applied to mental health support chatbots, our approach demonstrates that fine-tuned simulators dramatically outperform zero-shot base models at surfacing system issues, and adversarial training further enhances diversity, distributional alignment, and predictive validity. The resulting simulator achieves a strong correlation between simulated and real failure occurrence rates across diverse chatbot configurations while maintaining low distributional divergence of failure modes. Discriminator accuracy decreases drastically after three adversarial iterations, suggesting improved realism. These results provide evidence that adversarial training is a promising approach for creating realistic user simulators in mental health support TOD domains, enabling rapid, reliable, and cost-effective system evaluation before deployment.
Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization
@@ 1 Introduction
Task-oriented dialogue (TOD) systems guide users through multi-turn conversations to achieve specific goals (schatzmann2006survey). The development of large language model (LLM) based TOD systems has created an urgent need for robust evaluation methodologies that go beyond single-turn metrics. Multi-turn task-oriented conversations involve complex dynamics where system quality emerges from sustained interaction, user satisfaction accumulates over time, and task success depends on maintaining context across extended dialogue (ouyang2022training; hong2023zeroshotgoaldirecteddialoguerl). However, traditional evaluation through live user testing is expensive, time-consuming, and exposes users to potentially suboptimal system behavior.
User simulation offers a compelling alternative: synthetic agents that replicate human conversational behavior enable rapid iteration, systematic exploration of edge cases, and controlled experimentation (verma2022chai; yu2023promptbased). At the current state of the art, realistic user simulators have become essential for training language models through reinforcement learning (RL), where they serve as the environment for policy optimization (snell2023offlinerlnaturallanguage; zhou2024archertraininglanguagemodel). Recent advances in multi-turn dialogue optimization have demonstrated the potential of RL methods to improve long-term conversation quality (shani2024multiturnreinforcementlearningpreference; hong2024interactivedialogueagents). However, creating simulators that faithfully reproduce the nuanced, diverse, and goal-directed nature of human conversational behavior presents significant challenges. Simulators that fail to reproduce realistic user behavior cannot reliably expose system failure modes during development, leading to policies that appear effective in simulation but perform poorly with actual users.
In this work, we address the challenge of creating realistic user simulators through adversarial training. Inspired by Generative Adversarial Networks (GANs) (goodfellow2014generative), we train a discriminator to distinguish between simulated and real user sessions, then use the discriminator feedback to refine the simulator through Direct Preference Optimization (DPO) (rafailov2024direct). This creates a competitive dynamic where the simulator learns to fool an increasingly sophisticated discriminator, automatically identifying and correcting the most prominent simulation artifacts.
We evaluate our approach in the context of Ash [^1] https://www.talktoash.com/, a mental health support chatbot designed by Slingshot AI—a complex TOD domain where sustained conversation quality has direct implications for user wellbeing. Our experimental results demonstrate several key findings:
  • Fine-tuning dramatically outperforms zero-shot: Domain-specific fine-tuning is essential for capturing realistic user behavior distributions in TOD, including negative and resistant patterns that zero-shot models fail to exhibit.
@@
  • Adversarial training enhances realism: Three iterations of adversarial refinement progressively recover lexical diversity suppressed by supervised learning, increase distributional similarity, and dramatically reduce discriminator accuracy from near-perfect detection to near-random performance.
@@
  • Strong predictive validity: The best adversarial simulator achieves strong correlation with real model performance and successfully replicates A/B test results on conversation models, enabling reliable offline TOD evaluation before deployment.
@@
These contributions provide evidence that adversarial training is a promising methodology for creating realistic user simulators in the mental health support TOD domain, with potential applications to RL-based policy optimization and reliable offline evaluation. Future work will validate the approach across other TOD domains and integrate simulators into RL training loops.
@@ 2.1 Multi-Turn Dialogue Optimization
While RLHF (ouyang2022training) demonstrated that language models can be aligned with human preferences, it typically optimizes single-turn responses rather than multi-turn dialogue outcomes. DPO (rafailov2024direct) offers an alternative to RLHF by directly optimizing the policy without explicit reward modeling, simplifying the training pipeline while maintaining effectiveness.
Optimizing dialogue systems for long-term conversation quality requires methods that go beyond single-turn reward modeling. Recent approaches include training on imagined conversations and hindsight regeneration (hong2023zeroshotgoaldirecteddialoguerl; hong2024interactivedialogueagents), token-level policy optimization (shani2024multiturnreinforcementlearningpreference), hierarchical decomposition of planning and execution (zhou2024archertraininglanguagemodel), and offline RL methods for language generation (snell2023offlinerlnaturallanguage). These methods critically depend on the quality of the training environment, with realistic user simulators being essential for both policy optimization and evaluation (nam2025efficient).
@@ 2.2 User Simulation for Task-Oriented Dialogue
User simulation has been central to TOD research since early agenda-based systems that represented behavior through goal-directed dialogue acts (schatzmann2006survey; schatzmann2007agenda). Neural approaches later enabled more flexible simulation by capturing dialogue history without rigid structures (elasri2016usersim; kreyssig2018neural), demonstrating that policies trained on neural simulators could achieve higher success rates with real users. Recent work on mental health support benchmarking (pombal2024mindeval) demonstrates that LLMs struggle with longer interactions and patients presenting severe symptoms, underscoring the need for realistic simulators that capture the full complexity of human emotions and behaviors.
@@
LLM-Based Simulation: Large language models have transformed user simulation through in-context learning and prompting approaches (li2022dialogic; davidson2023prompting), achieving goal success rates comparable to human-bot interactions with greater linguistic diversity and minimal seed dialogues. However, prompting-only approaches suffer from hallucinations and goal drift. Fine-tuning addresses these limitations through various techniques, including reducing behavioral inconsistencies, joint policy and NLG optimization for domain transfer, and dual-model architectures with chain-of-thought reasoning (sekulic2024daus; lin2022gentus; luo2024duetsim).
RL-Based Simulator Training: While supervised learning dominates simulator training, reinforcement learning approaches optimize for task-level objectives rather than imitation alone. Multi-agent RL enables joint system-user training (takanobu2020guided; liu2018dialogue), while adversarial training creates competitive dynamics between simulators and discriminators (keizer2023adversarial; li2017adversarial). The adversarial approach has demonstrated substantial improvements, with discriminator-based rewards being "more general, natural and effective" than hand-crafted metrics. Our work extends this adversarial paradigm to LLM-based simulators, using DPO rather than policy gradient methods to leverage discriminator feedback while avoiding reward hacking.
@@
Evaluation Methodology: Traditional TOD evaluation relies on task-level metrics such as success rate and slot error rate measured on standard benchmarks like MultiWOZ (budzianowski2018multiwoz), SGD (rastogi2020schema), and tau-bench (yao2024taubench), which effectively capture transactional outcomes but provide limited insight into conversational quality. The poor correlation between surface-level metrics like BLEU/ROUGE and human judgments (liu2016not) motivates more sophisticated evaluation approaches that consider behavioral fidelity and distributional alignment (nam2025efficient). Our work advances this evaluation paradigm through three key contributions: (1) explicit modeling of conversational feature occurrence rate distributions using KL divergence to measure alignment on fine-grained failure modes rather than binary task success, (2) validation of predictive utility through model A/B test replication beyond correlation analysis, and (3) discriminator accuracy as a direct measure of distributional realism.
@@ 3.2 Context Construction for User Simulation
A critical design choice for realistic user simulation is the construction of context $c$ that conditions the simulator’s behavior. Effective context should provide: (1) role and behavioral guidelines through system instructions, (2) user background and goals, (3) conversational history to maintain consistency, and (4) session-specific grounding to anchor the simulator to target conversation characteristics. The specific context structure depends on the domain—transactional TOD systems might include task schemas and slot values, while open-ended domains might emphasize user profiles and session summaries. Our mental health support chatbot implementation (detailed in Appendix A) uses a hierarchical structure spanning user profiles, multi-session summaries, and current session grounding. As we demonstrate in Section 5, rich context grounding enables fine-tuned simulators to cluster tightly with their corresponding real sessions in embedding space, while zero-shot models without such grounding form distinct clusters, indicating systematic behavioral drift.
@@ 3.3 Base Simulator Training
We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations. Starting from Llama-3.3-70B-Instruct (grattafiori2024llama3) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
@@ 3.4 Discriminator Training
To measure simulator realism, we train a discriminator $D_{\phi}$ to distinguish between simulated and real user sessions. We implement the discriminator as a token classification model built on the Llama-3.1-8B-Instruct backbone. The discriminator processes the full conversation and predicts after each user message whether the session is real or simulated:
@@
We use causal attention in the discriminator, meaning predictions are made based on the current message and all previous messages in the session. This design choice is essential because any individual message could plausibly be either real or simulated when taken out of context. The discriminator’s effectiveness stems from detecting distributional patterns across the conversation: certain phrases repeated multiple times, unnatural topic progressions, or behavioral patterns that diverge from real user dynamics. Without conversational context, the discrimination task becomes noisy and ill-defined, as isolated messages lack the sequential dependencies that reveal simulation artifacts.
@@
In our adversarial training approach, each new discriminator is trained exclusively on data from the most recent simulator version. This ensures the preference pairs we generate for DPO reflect the current policy’s behavior rather than artifacts from earlier iterations. Research on preference optimization supports this on-policy approach: on-policy preference pairs exhibit stronger contrastiveness and clearer preference signals (zhao2024reflective), online methods require only “partial coverage” compared to the strict “global coverage” needed by offline methods (song2024importance), and on-policy sampling efficiently concentrates probability mass on high-reward responses (tajwar2024preference).
3.5 Adversarial Training Framework
Our adversarial training process follows an iterative refinement loop, illustrated in Figure 1. Algorithm 1 provides the detailed procedure.
Algorithm 1 Adversarial User Simulator Training
@@ 3.5 Adversarial Training Framework
Figure 1: Overview of the adversarial training process. The simulator generates sessions that are evaluated by the discriminator to distinguish real from simulated conversations. Discriminator feedback guides preference dataset generation, which is then used to refine the simulator through DPO, creating an iterative competitive dynamic that improves realism.
Optimization Method: We use DPO rather than RLHF methods (schulman2017proximal; shao2024deepseekmath) because discriminator rewards are inherently policy-dependent and vulnerable to reward hacking. With RLHF, the simulator would continuously optimize against a fixed reward model, eventually exploiting patterns that score highly but don’t reflect genuine realism. Preventing this would require retraining the discriminator throughout optimization—a substantial computational overhead. DPO avoids this problem: we generate preference pairs once per iteration using the current discriminator, then optimize against this fixed dataset. Preferences grounded in comparisons between specific alternatives are more robust than absolute reward values, and the discrete iteration structure naturally prevents reward hacking.
@@
Preference Dataset Generation: At messages with the highest and lowest rewards (critical moments where the discriminator most clearly detects unrealism), we sample 8 alternative responses. These alternatives are ranked by their discriminator-based rewards to create preference pairs $(u_{\text{chosen}},u_{\text{rejected}})$ where the chosen response has the highest reward.
@@
DPO Update: We train the new simulator using Direct Preference Optimization (rafailov2024direct), which directly optimizes the policy to prefer higher-reward responses. We use the previous iteration’s simulator as the reference policy $\pi_{\text{ref}}$. Defining the implicit reward $r_{\theta}(h,c,u)=\log\frac{\pi_{\theta}(u|h,c)}{\pi_{\text{ref}}(u|h,c)}$, the DPO loss is:
@@ 4.1 Domain and Data
We evaluate our methodology in the context of mental health support—a complex task-oriented dialogue domain where users pursue mental health goals (emotional regulation, insight development, behavioral change) through extended multi-turn conversations. Unlike traditional TOD domains with discrete transactional goals, mental health support offers distinct advantages for studying realistic user simulation: rich emotional dynamics and behavioral complexity, multi-session continuity, diverse naturalistic failure modes beyond binary task success, and production-scale data capturing rare but critical patterns like crisis expressions and resistance to engagement.
Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).

### v3

# DIAL: Direct Iterative Adversarial Learning for Realistic Multi-Turn Dialogue Simulation

@@ Abstract
Realistic user simulation is crucial for training and evaluating multi-turn dialogue systems, yet creating simulators that accurately replicate human behavior remains a significant challenge. An effective simulator must expose the failure modes of the systems under evaluation. This work introduces Direct Iterative Adversarial Learning (DIAL), a DPO-based adversarial training framework that iteratively enhances user simulator realism through a competitive dynamic between a generator (user simulator) and a discriminator. When applied to mental health support, a domain characterized by diverse failure types and a critical dependence on realistic user behavior for failure detection, DIAL restores lexical diversity diminished by supervised fine-tuning and reduces discriminator accuracy from near-perfect to near-random levels. The resulting simulator exhibits a strong correlation between simulated and real failure occurrence rates while maintaining low distributional divergence of failure modes. These findings indicate that DIAL is a promising method for developing realistic user simulators in multi-turn dialogue, facilitating rapid, reliable, and cost-effective system evaluation prior to deployment.
DIAL: Direct Iterative Adversarial Learning for Realistic Multi-Turn Dialogue Simulation
@@ 1 Introduction
Task-oriented dialogue (TOD) systems facilitate multi-turn conversations to help users achieve specific objectives (Schatzmann et al., 2006). The emergence of large language models (LLMs) has highlighted the need for robust evaluation methodologies that extend beyond single-turn metrics, as system quality emerges from sustained interaction and task success depends on maintaining context across extended dialogue. Traditional evaluation methods, such as live user testing, are costly, time-intensive, and may expose users to suboptimal system behavior.
User simulation provides a viable alternative, as synthetic agents that mimic human conversational behavior enable rapid iteration, systematic exploration of edge cases, and controlled experimentation (Verma et al., 2022; Yu et al., 2023). However, developing simulators that accurately capture the nuanced, diverse, and goal-oriented nature of human conversation remains challenging. Simulators that do not reflect realistic user behavior fail to reveal system failure modes during development, resulting in policies that perform well in simulation but poorly with real users.
This work addresses the challenge of developing realistic user simulators through Direct Iterative Adversarial Learning (DIAL). Drawing inspiration from Generative Adversarial Networks (GANs) (Goodfellow et al., 2014), DIAL trains a discriminator to differentiate between simulated and real user sessions, and uses discriminator feedback to refine the simulator via Direct Preference Optimization (DPO) (Rafailov et al., 2023). Unlike prior adversarial dialogue approaches that rely on policy gradient methods (Li et al., 2017; Zhao et al., 2021), DIAL avoids training instability and naturally prevents reward hacking through its discrete iteration structure (see Section 3 for details).
DIAL is evaluated using Ash [^1] https://www.talktoash.com/, a mental health support chatbot developed by Slingshot AI. The mental health domain is selected due to its diverse range of failure types, which require realistic user simulation for effective detection of system shortcomings. Experimental results reveal several key findings:
@@
  • DIAL methodology: Three iterations of DIAL refinement using DPO on discriminator-derived preference pairs progressively restore lexical diversity diminished by supervised learning, enhance distributional similarity, and reduce discriminator accuracy from near-perfect to near-random levels.
  • Failure-sensitive evaluation: Domain-specific fine-tuning, in conjunction with DIAL refinement, is critical for capturing realistic user behaviors, including negative and resistant patterns that zero-shot models do not exhibit. This enables accurate identification of diverse chatbot failure modes.
@@
  • Strong predictive validity: The optimal DIAL-trained simulator demonstrates a strong correlation with real model performance and accurately replicates A/B test results on conversation models, facilitating reliable offline evaluation.
@@
These findings indicate that DIAL is a promising methodology for developing realistic user simulators in mental health support dialogues, with potential applications in RL-based policy optimization and reliable offline evaluation. Future research will extend validation to additional TOD domains and incorporate simulators into end-to-end training processes.
@@ 2.1 Multi-Turn Dialogue Optimization
Although RLHF (Ouyang et al., 2022) has shown that language models can be aligned with human preferences, its optimization is generally limited to single-turn responses. DPO (Rafailov et al., 2023) provides an alternative by directly optimizing the policy without explicit reward modeling, thereby simplifying the training pipeline while preserving effectiveness.
Enhancing dialogue systems for sustained conversational quality necessitates approaches that extend beyond single-turn reward modeling. Recent methods encompass training on imagined conversations and hindsight regeneration (Hong et al., 2023, 2024), token-level policy optimization (Shani et al., 2024), hierarchical decomposition of planning and execution (Zhou et al., 2024), and offline reinforcement learning for language generation (Snell et al., 2023). The effectiveness of these approaches is highly contingent on the quality of the training environment, with realistic user simulators being indispensable for both policy optimization and evaluation (Nam et al., 2025).
@@ 2.2 User Simulation for Task-Oriented Dialogue
User simulation has played a central role in TOD research since the development of early agenda-based systems, which modeled behavior through goal-directed dialogue acts (Schatzmann et al., 2006, 2009). Subsequent neural approaches enabled more flexible simulation by capturing dialogue history without rigid structural constraints (El Asri et al., 2016; Kreyssig et al., 2018), and demonstrated that policies trained on neural simulators can achieve higher success rates with real users. Recent benchmarking in mental health support (Pombal et al., 2025) reveals that LLMs encounter difficulties with extended interactions and patients exhibiting severe symptoms, highlighting the necessity for realistic simulators that reflect the full complexity of human emotions and behaviors.
@@
LLM-Based Simulation: LLMs have significantly advanced user simulation by leveraging in-context learning and prompting methods (Li et al., 2022; Davidson et al., 2024), achieving goal success rates comparable to human-bot interactions while offering greater linguistic diversity and requiring minimal seed dialogues. Nevertheless, prompting-only strategies are prone to hallucinations and goal drift. Fine-tuning mitigates these issues through techniques such as reducing behavioral inconsistencies, jointly optimizing policy and natural language generation (NLG) for domain transfer, and employing dual-model architectures with chain-of-thought reasoning (Sekulić et al., 2024; Lin et al., 2022; Luo et al., 2024).
RL-Based Simulator Training: While supervised learning dominates simulator training, reinforcement learning methods optimize for task-level objectives beyond mere imitation. Multi-agent reinforcement learning facilitates joint system-user training (Takanobu et al., 2020; Liu and Lane, 2017), while adversarial training introduces competitive dynamics between simulators and discriminators (Keizer et al., 2023; Li et al., 2017; Zhao et al., 2021). Li et al. (2017) employ REINFORCE with Monte Carlo search to backpropagate discriminator rewards, but report that this leads to unstable training and necessitates teacher forcing as a regularizer. Zhao et al. (2021) also utilize GAN-based policy gradient methods for user simulation in recommendation systems. The present work contributes methodologically by extending the adversarial paradigm to LLM-based simulators through DIAL, with a key distinction in optimization: DPO is applied to discriminator-derived preference pairs instead of policy gradient methods. DIAL circumvents the instability associated with policy gradients in text generation and inherently prevents reward hacking (see Section 3).
@@
Evaluation Methodology: Traditional TOD evaluation, including recent end-to-end frameworks such as AutoTOD (Xu et al., 2024) and ProTOD (Dong et al., 2025), relies on task-level metrics such as success rate and slot accuracy measured on standard benchmarks like MultiWOZ (Budzianowski et al., 2018), SGD (Rastogi et al., 2020), and tau-bench (Yao et al., 2024). These metrics effectively capture transactional outcomes but provide limited insight into conversational quality. Our work addresses the complementary problem of building realistic user simulators to evaluate such systems, focusing on failure mode detection rather than transactional success, and advances this evaluation paradigm through three key contributions: (1) explicit modeling of conversational feature occurrence rate distributions using KL divergence to measure alignment on fine-grained failure modes, (2) validation of predictive utility through model A/B test replication beyond correlation analysis, and (3) discriminator accuracy as a direct measure of distributional realism.
@@ 3.2 Context Construction for User Simulation
A critical design choice for realistic user simulation is the construction of context $c$ that conditions the simulator’s behavior. Effective context should provide: (1) role and behavioral guidelines, (2) user background and goals, (3) conversational history to maintain consistency, and (4) task-specific grounding to anchor the simulation. The specific context structure depends on the domain—transactional TOD systems might include task schemas and slot values, while open-ended domains might emphasize user profiles and session summaries. Our mental health support chatbot (detailed in Appendix A) uses a hierarchical structure spanning user profiles, multi-session summaries, and current session grounding. As we demonstrate in Section 5, rich context grounding enables fine-tuned simulators to cluster tightly with their corresponding real sessions in embedding space, while zero-shot models exhibit systematic behavioral drift.
@@ 3.3 Base Simulator Training
We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations. Starting from Llama-3.3-70B-Instruct (Grattafiori et al., 2024) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
@@ 3.4 Discriminator Training
To measure simulator realism, we train a discriminator $D_{\phi}$ to distinguish between simulated and real user sessions. We implement the discriminator as a token classification model which processes the full conversation and predicts after each user message whether the session is real or simulated:
@@
We use causal attention in the discriminator, meaning predictions are made based on the current message and all previous messages in the session. The discriminator’s effectiveness stems from detecting distributional patterns across the conversation: repeated phrases, unnatural topic progressions, or behavioral patterns that diverge from real user dynamics. Without conversational context, the discrimination task becomes noisy and ill-defined, as isolated messages lack the sequential dependencies that reveal simulation artifacts.
@@
In DIAL, each new discriminator is trained exclusively on data from the most recent simulator version. This ensures the preference pairs we generate for DPO reflect the current policy’s behavior rather than artifacts from earlier iterations. Research on preference optimization supports this on-policy approach: on-policy preference pairs exhibit stronger contrastiveness and clearer preference signals (Zhao and Li, 2025), online methods require only “partial coverage” compared to the strict “global coverage” needed by offline methods (Song et al., 2024), and on-policy sampling efficiently concentrates probability mass on high-reward responses (Tajwar et al., 2024).
3.5 DIAL: Direct Iterative Adversarial Learning
The DIAL training process follows an iterative refinement loop, illustrated in Figure 1. Algorithm 1 provides the detailed procedure.
Algorithm 1 DIAL: Direct Iterative Adversarial Learning
@@ 3.5 DIAL: Direct Iterative Adversarial Learning
Figure 1: Overview of the DIAL training process. The simulator generates sessions that are evaluated by the discriminator to distinguish real from simulated conversations. Discriminator feedback guides preference dataset generation, which is then used to refine the simulator through DPO, creating an iterative competitive dynamic that improves realism.
Optimization Method: DIAL uses DPO rather than RLHF methods (Schulman et al., 2017; Shao et al., 2024) because discriminator rewards are inherently policy-dependent and vulnerable to reward hacking. With RLHF, the simulator would continuously optimize against a fixed reward model, eventually exploiting patterns that score highly but don’t reflect genuine realism. Preventing this would require retraining the discriminator throughout optimization—a substantial computational overhead. DPO avoids this problem: we generate preference pairs once per iteration using the current discriminator, then optimize against this fixed dataset. Preferences grounded in comparisons between specific alternatives are more robust than absolute reward values, and the discrete iteration structure naturally prevents reward hacking.
@@
Preference Dataset Generation: At messages with the highest and lowest rewards, we sample 8 alternative responses. These alternatives are ranked by their discriminator-based rewards to create preference pairs $(u_{\text{chosen}},u_{\text{rejected}})$ where the chosen response has the highest reward.
@@
DPO Update: We train the new simulator using Direct Preference Optimization (Rafailov et al., 2023), which directly optimizes the policy to prefer higher-reward responses. We use the previous iteration’s simulator as the reference policy $\pi_{\text{ref}}$. Defining the implicit reward $r_{\theta}(h,c,u)=\log\frac{\pi_{\theta}(u|h,c)}{\pi_{\text{ref}}(u|h,c)}$, the DPO loss is:
@@ 4.1 Domain and Data
DIAL is evaluated within the domain of mental health support, in which users pursue mental health goals such as emotional regulation, insight development, and behavioral change through extended multi-turn conversations. This domain was selected to specifically assess simulators’ capacity to reveal failure modes of TOD systems, as mental health conversations present a particularly diverse array of failure types (see Appendix E). In contrast to traditional TOD benchmarks such as MultiWOZ (Budzianowski et al., 2018) and SGD (Rastogi et al., 2020), where failure modes are primarily transactional (e.g., incorrect slot filling or task incompletion) and less reliant on realistic user simulation, the mental health support domain features complex emotional dynamics and naturalistic failure modes that require faithful user behavior for effective detection. Furthermore, production-scale data in this domain captures rare but critical phenomena, such as crisis expressions and resistance to engagement, thereby providing a rigorous testbed for evaluating simulator realism.
Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context.